OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

OpenStudy (anonymous):

@divu.mkr

OpenStudy (anonymous):

take 2 common out of mod as it will not affect the mod.. than take x<-3 -(x+3)=2 then x>-3 x+3=2

OpenStudy (anonymous):

-9 on the first? 15?

OpenStudy (jdoe0001):

\(\bf -3|2x + 6| = -12 \implies |2x + 6| = \cfrac{-12}{-3} \implies \begin{cases} +1(2x + 6) = \cfrac{-12}{-3}\\\quad \\ \bf -1(2x + 6) = \cfrac{-12}{-3} \end{cases}\)

OpenStudy (anonymous):

no -5 nd -1

OpenStudy (anonymous):

sorry i was looking at the wrong problem..

OpenStudy (anonymous):

no worrries (: